|Deletions are marked like this.||Additions are marked like this.|
|Line 36:||Line 36:|
|Date: 2022/8/19 14:00-||Date: 2022/8/19 14:00-14:15|
|Line 38:||Line 38:|
|Participants:||Participants: Takahashi, Fabian, Tanaka, Aso|
CRY+VIS Interface Meeting
Date: 2022/8/19 13:00-13:20
Participants: Takahashi, Uchiyama, Yokozawa, Miyoki, Tomaru, Ushiba, Akutsu, Hirata, Yasui, Ikeda, Node
- Type-A: Offload of BF GAS in IY.
Type-B: Polarity Check of magnets. Investigation of F1 damper. Recovery of geophone signals in SR2/SRM. Preparation for Review.
- Type-Bp: Oplev rearrangement in PR3. Taking 360deg pictures inside the chambers in the next week.
- Replacement of cables: EXT and IY/EX remain.
- TM height setting: Type-B in the next week.
- Oplev: Check of the beam pass in the next week.
- Commissioning in air:
- Commissioning in vacuum: IR beam came to the AS port.
Acceptance check: DropboxPaper
- New vacuum pumps: Side hatch in SRM on 8/23.
- Heater control: Pending.
- TMS: Replacement of pressure gauge in EX.
Date: 2022/8/19 14:00-14:15
Participants: Takahashi, Fabian, Tanaka, Aso
- Offloaded the BF GAS in IY.
I posted plots comparing BS and SR2 TFs calculated with the Matlab rigid-body model. See klog 21813
- Review at 15:00 on 19th.
- Type-B: Height check. Rubbing check. Adjustment of F1 damper in BS/SR2. Adjustment of F0 FR in BS. Oplev check (various).
- How to set the F0 FR in BS?
The present temperature is 23.5. Though we can use the heater on the chamber later, we should keep a margine of the cooling power. 23.2 is optimal?
(Date in Kamioka)
The week from 8/22
- Takahashi: 8/22-26
- Sato: 8/22-26
- Hirata: 8/22-26
The week from 8/29